# Towards Building Multilingual Language Model for Medicine

### v2

@@ 4.2 Auto-regressive Training on MMedC
Once constructed MMedC, we further pre-train the existing LLMs on it in an auto-regressive manner. We adopt loss for next token prediction as used in GPT [20]. Specifically, we treat medical text as a sequence of tokens, denoted as $X=\{x_{1},x_{2},\dots,x_{N}\}$, where each $x_{i}$ is a text token and $N$ represents the total length of the sequence. For a token $x_{i}$ to be predicted, the optimization objective is:
@@ 4.3 Comprehensive Multilingual Medical Benchmark
In addition to the multilingual datasets for training, we also collect a comprehensive Multilingual Medical Benchmark, spanning 6 principal languages, namely MMedBench, to conduct a thorough evaluation of a model’s performance in the medical domain across diverse languages Specifically, we started by collecting existing medical question-answering (QA) benchmarks for each language, and expand these multi-choice QA with corresponding explanations using GPT-4, followed by strict human verification to ensure the correctness of contents.
4.3.1 Multilingual Medical QA Dataset
Figure 5: The pipeline of MMedB construction. Firstly multi-choice QA pairs from various languages are collected from 5 QA datasets. Then corresponding rationale is generated with the help of GPT4. The rationale of testset is further checked by humans to ensure its quality.
Evaluating the performance of Large Language Models (LLMs) has conventionally relied on the utilization of multiple-choice questions. This evaluation framework presents the question and its corresponding options to the model, which is then expected to identify the correct answer’s index. Accuracy serves as the primary quantitative metric in this method, providing a direct and objective measure of the performance. Despite its efficacy, the prevalent medical multi-choice QA benchmarks are exclusively monolingual, thus falling short of adequately assessing LLMs’ capabilities across diverse languages. To address this deficiency and foster a more inclusive evaluation landscape, our approach involves the aggregation of various medical multi-choice QA datasets from multiple languages. This initiative aims to compile a comprehensive benchmark that reflects the multilingual realities of the medical field. The following benchmarks are considered:
@@ 4.3.1 Multilingual Medical QA Dataset
4.3.2 Rationale Generation
While the accuracy of multi-choice question answering is a straightforward and accurate metric, it fails to evaluate the reasoning and long sentence generation abilities of LLMs, which is critical for clinical usage. Therefore, we further augment each question with a justification for selecting the correct option. At evaluation phase, we prompt the model to articulate the rationale behind its choice, thereby offering insights into the model’s reasoning capabilities, as shown by Figure 5.
@@ 4.3.2 Rationale Generation
You’re a {language} doctor. Analyze the reasons behind choosing this particular option in 100 words for the following question in {language}.
@@
Following the rationale generation by GPT-4, we conducted manual review to assess their quality. Our evaluation criteria are two-fold: firstly, the explanation provided by GPT-4 had to be consistent with the established correct answer for the question; secondly, it’s required to articulate the logic underpinning the answer, rather than merely replicating it. Note that, considering the cost of human checks, this will only be performed on part of our test set. Specifically, we randomly selected 200 samples from the former test split for each language to form a new rationale split. Then we shuffled and distributed among three annotators for manual verification. Annotators were tasked with categorizing each rationale as either qualified or unqualified based on the aforementioned criteria. Remarkably, we observed that 94.7% of the rationales generated by GPT-4 adhered to our standards, underscoring the high quality of the explanations. During the final evaluation phase, the calculation of rationale similarity was exclusively applied to these human-verified passed samples.
Finally, we get 1136 human-checked samples for rationale evaluation and compensate the former split 45048 training QA pairs with auto-generated rationale sentences. Given a language model, it can use our training set to further fine-tune and then be evaluated or directly evaluated on the rationale and choice testing sets.
4.3.3 Topic classification
@@
Here, we explore the thematic distribution of the samples. For this purpose, we employ GPT-4 to categorize the topics within the test set. The instructions provided to GPT-4 for topic classification are outlined as follows, where similarly, “{language}” will be replaced by a certain language name.: You’re a {language} doctor, choose one subject out of {medical_subjects_string}, which is most relevant to the following question. At times, GPT-4 may yield ambiguous classification outcomes, such as those not aligning with predefined medical subjects. In these instances, we prompt GPT-4 to try the classification up to four times. Should it fail to produce a categorization that adheres to our criteria, we assign the sample’s category as ‘None’. Given the rarity of such occurrences, their impact on the overall statistics is minimal.
4.3.4 Evaluation Settings
@@
To comprehensively assess the model’s performance, we tested it in three different evaluation settings: zero-shot, parameter efficient fine-tuning (PEFT), and full fine-tuning. For the zero-shot setting, we directly test off-the-shelf LLMs with proper instruction, without any further exposure to the training part of MMedBench. In addition to zero-shot, to better evaluate the performance differences between models, we also try to fine-tune and then test the open-source models. There are two mainly used fine-tuning approaches: parameter efficient fine-tuning (PEFT), and full fine-tuning. In former case, only a small subset of the model’s parameters are trainable, representing the performance in low computing resource available scenario. We adopt the most representative PEFT method LoRA [7] here. In latter case, all parameters will be fine-tuned which is a more conventional practice.
We have benchmarked the following LLMs on our proposed MMedBench:
@@
  • GPT-4 [20], a groundbreaking multilingual large language model developed by OpenAI, stands as one of the most sophisticated LLMs to date. Due to the confidential of data and model details, it is uncertain for its detailed model scale. Though GPT-4 does not emphasize it is a multilingual LLM, its multilingual ability is still superior. Given that it is only accessible through API, we evaluate it in a zero-shot manner with proper instruction (Check Supplementary Material A for more details).
@@
  • Gemini-1.0 pro [24] is the latest general multimodal foundation model developed by Google. Though it is targeted at multimodal scenarios, as reported in the original paper, its language ability even surpasses Google’s former LLM, PaLM 2 [2]. Similar to GPT-series, its detailed scale and whether it is specifically targeted to multilingual or monolingual scenarios are not released. However, in our testing, it responds well for multilingual input.
@@
  • PMC-LLaMA [29] presents another open-source monolingual medical-specific LLM, pretrained exclusively on English medical literature, including papers and books. We utilize the 7B parameter version for evaluation.
  • LLaMA 2 [26], the latest iteration in the LLaMA series, is recognized as one of the most potent open-source monolingual LLMs with a focus on English. We engage the 7B parameter model in our finetuning evaluation process.
@@
4.3.5 Metrics and Human Rating
To evaluate the performance of LLMs, we employ two metrics: Accuracy and Rationale Similarity. Measuring Accuracy is straightforward, as the LLM can generate outputs following a specific template. However, assessing Rationale Similarity presents a more complex challenge, typical within the NLP domain. Initially, we applied three classical text similarity methods, BLEU [21] and ROUGE [16], and Bert-score [34].
@@ 4.3.5 Metrics and Human Rating
BERT-score utilizes the contextual embedding from a pretrained BERT to capture high-level semantic features, calculating the similarity between reference and candidate texts through cosine similarity. The Recall is calculated as follows:
@@
Relative Rating scores aim to rank the output based on relative comparison. In detail, we select 6 representative models and sample 50 cases for each language. In human rating, for each case, question, options, right answer and rationale generated by each models are present to annotators, along with the reference rationale. The annotators are asked to rank the model-generated rationales based on the following three evaluation criteria:
@@ 4.4 Implementation Details
Auto-regressive Training. During further auto-regressive training on MMedC, our optimization objective aligns with that of the auto-regressive generation task. For data processing, we segment the text into chunks, each comprising 2048 tokens, with an overlapping margin of 512 tokens. Throughout the training, we maintain a maximum context length of 2048 tokens. Owing to the model’s extensive parameter count, which precludes fitting on a single GPU, we employ the Fully Sharded Data Parallel (FSDP) strategy to distribute the model across multiple GPU cards. Additionally, we utilize the BF16 data type and gradient checkpointing techniques to optimize memory usage. For InternLM, we establish a global batch size of 512 and a learning rate of 2e-5. In the case of BLOOM, we set the global batch size to 512 and a learning rate of 8e-6. We pretrain both models on eight A100 GPUs, adapting gradient accumulation steps to sustain such a large global batch size. We train the 7B model for 20k iterations, which takes about 20 Days.
@@ 5 Conclusion
In this paper, we describe an automatic pipeline for building up a new multilingual medical corpus together with a new benchmark. Specifically, we developed MMedC, a large-scale medical corpus with 25.5B tokens covering six main languages, and MMedBench, a comprehensive benchmark encompassing 6 primary languages. To monitor the development progress of medical multilingual LLMs, we have evaluated eleven existing LLMs on multi-choice question-answering and rationale generation abilities under various settings. Experimentally, we demonstrate the effectiveness of further training on MMedC, significantly filling the gap in adapting advanced general multilingual LLMs into complex and professional medical domain. As a result, we open-source the MMedLM 2, which to our knowledge, denotes the first, and strongest multilingual language model for medicine.
6 Data and Code Availability
Currently, we have released all of our data, code, and model weights.
@@ A.0.1 Zero-shot prompt
For models such as GPT-3.5 and GPT-4, we employ specific instructions to steer the model’s response to questions. To prompt the model for a straightforward correct answer, we use the following directive: You’re a {language} doctor, make a choice based on the question and options. You need to answer the letter of the option instead of answering the entire option or anything else. Options may not be unique. When we aim for the model to provide not just the correct answer but also its corresponding rationale, we utilize a different instruction: You’re a {language} doctor, make a choice based on the question and options in {language}. You should solve this step-by-step. You must first give the reason in {language} for your choice ends with ’[End]’. Then you must give the answer’s letter directly again. The template is like ’Reason:… [End] Answer: A, B’

### v3

@@ 4.2 Auto-regressive Training on MMedC
Once constructed MMedC, we further pre-train the existing LLMs on it in an auto-regressive manner. We adopt loss for the next token prediction as used in GPT [1]. Specifically, we treat medical text as a sequence of tokens, denoted as $X=\{x_{1},x_{2},\dots,x_{N}\}$, where each $x_{i}$ is a text token and $N$ represents the total length of the sequence. For a token $x_{i}$ to be predicted, the optimization objective is:
@@ 4.3 Comprehensive Multilingual Medical Benchmark
In addition to the multilingual datasets for training, we also collect a comprehensive Multilingual Medical Benchmark, spanning 6 principal languages, namely MMedBench, to conduct a thorough evaluation of a model’s performance in the medical domain across diverse languages. Specifically, we started by collecting existing medical question-answering (QA) benchmarks for each language, and expand these multi-choice QA with corresponding explanations using GPT-4, followed by strict human verification to ensure the correctness of contents.
Figure 5: The pipeline of MMedBench construction. Firstly multi-choice QA pairs from various languages are collected from 5 QA datasets. Then corresponding rationale is generated with the help of GPT4. The rationale of testset is further checked by humans to ensure its quality.
Multilingual Medical QA Dataset. Evaluating the performance of Large Language Models (LLMs) has conventionally relied on the utilization of multiple-choice questions. This evaluation framework presents the question and its corresponding options to the model, which is then expected to identify the correct answer’s index. Accuracy serves as the primary quantitative metric in this method, providing a direct and objective measure of the performance. Despite its efficacy, the prevalent medical multi-choice QA benchmarks are exclusively monolingual, thus falling short of adequately assessing LLMs’ capabilities across diverse languages. To address this deficiency and foster a more inclusive evaluation landscape, our approach involves the aggregation of various medical multi-choice QA datasets from multiple languages. This initiative aims to compile a comprehensive benchmark that reflects the multilingual realities of the medical field. The following benchmarks are considered:
@@
Rationale Generation. While the accuracy of multi-choice question answering is a straightforward and accurate metric, it fails to evaluate the reasoning and long sentence generation abilities of LLMs, which is critical for clinical usage. Therefore, we further augment each question with a justification for selecting the correct option. At evaluation phase, we prompt the model to articulate the rationale behind its choice, thereby offering insights into the model’s reasoning capabilities, as shown by Figure 5.
@@
[backgroundcolor=gray!20] You’re a {language} doctor. Analyze the reasons for choosing this particular option in 100 words for the following question in {language}.
@@
Following the rationale generation by GPT-4, we conducted manual review to assess their quality. Our evaluation criteria are two-fold: firstly, the explanation provided by GPT-4 had to be consistent with the established correct answer for the question; secondly, it’s required to articulate the logic underpinning the answer, rather than merely replicating it. Note that, considering the cost of human checks, this will only be performed on part of our test set. Specifically, we randomly selected 200 samples from the former test split for each language to form a new rationale split. Then we shuffled and distributed among three annotators for manual verification. Annotators were tasked with categorizing each rationale as either qualified or unqualified based on the aforementioned criteria. Remarkably, we observed that 94.7% of the rationales generated by GPT-4 adhered to our standards, underscoring the high quality of the explanations. During the final evaluation phase, the calculation of rationale similarity was exclusively applied to these human-verified passed samples. Finally, we get 1136 human-checked samples for rationale evaluation and compensate the former split 45048 training QA pairs with auto-generated rationale sentences. Given a language model, it can use our training set to further fine-tune and then be evaluated or directly evaluated on the rationale and choice testing sets.
@@
Topic Classification. Subsequently, we explore the thematic distribution of the samples. For this purpose, we employ GPT-4 to categorize the topics within the test set. The instructions provided to GPT-4 for topic classification are outlined as follows, where similarly, “{language}” will be replaced by a certain language name.: {mdframed}[backgroundcolor=gray!20] You’re a {language} doctor, choose one subject out of {medical_subjects_string}, which is most relevant to the following question. At times, GPT-4 may yield ambiguous classification outcomes, such as those not aligning with predefined medical subjects. In these instances, we prompt GPT-4 to try the classification up to four times. Should it fail to produce a categorization that adheres to our criteria, we assign the sample’s category as ‘None’. Given the rarity of such occurrences, their impact on the overall statistics is minimal.
@@
Evaluation Settings. To comprehensively assess the model’s performance, we tested it in three different evaluation settings: zero-shot, parameter efficient fine-tuning (PEFT), and full fine-tuning. For the zero-shot setting, we directly test off-the-shelf LLMs with proper instruction, without any further exposure to the training part of MMedBench. In addition to zero-shot, to better evaluate the performance differences between models, we also try to fine-tune and then test the open-source models. There are two mainly used fine-tuning approaches: parameter efficient fine-tuning (PEFT), and full fine-tuning. In the former case, only a small subset of the model’s parameters are trainable, representing the performance in a low computing resource available scenario. We adopt the most representative PEFT method LoRA [13] here. In the latter case, all parameters will be fine-tuned which is a more conventional practice.
Next, we will introduce the baseline LLMs considered in our work for comparison:
@@
  • GPT-4 [1], a groundbreaking multilingual large language model developed by OpenAI, stands as one of the most sophisticated LLMs to date. Due to the confidentiality of data and model details, it is uncertain for its detailed model scale. Though GPT-4 does not emphasize it is a multilingual LLM, its multilingual ability is still superior. Given that it is only accessible through API, we evaluate it in a zero-shot manner with proper instruction (Check Supplementary Material A for more details).
  • GPT-4 (5-shot, CoT) uses in context learning [5] and chain-of-thought [20] to further improve the performance of GPT-4, which represents the highest performance currently achievable. For implementation, we follow prompts used in MedPrompt[31]. Notice that, despite this approach can enhance the performance of different LLMs, it will take up more tokens and result in additional costs.
@@
  • Flan-PaLM [38] and MedPaLM 2 [39], is two close-source multilingual biomedical LLMs developed by Google. They demonstrate strong performance in medical English multiple-choice question answering. However, since it provides neither model weights nor access API function, we can only compare with them on the widely-used English benchmarks. A famous variant of Flan-PaLM is MedPaLM [38], while, in the original paper, the multiple-choice question answering accuracy of MedPaLM is not reported. Thus, here we can only compare with Flan-PaLM.
  • Gemini-1.0 pro [40] is the latest general multimodal foundation model developed by Google. Though it is targeted at multimodal scenarios, as reported in the original paper, its language ability even surpasses Google’s former LLM, PaLM 2 [2]. Similar to GPT series, its detailed scale and whether it is specifically targeted to multilingual or monolingual scenarios are not released. However, in our testing, it responds well for multilingual input.
@@
  • PMC-LLaMA [48] presents another open-source monolingual medical-specific LLM, pre-trained exclusively on English medical literature, including papers and books. We utilize the 7B parameter version for evaluation.
  • Llama 2 and Llama 3 [43], Llama series is a series of open-source LLMs developed by Meta. Llama 2 is the previous generation LLM in the series and Llama 3 is the latest one. Llama models are acknowledged as among the most powerful open-source monolingual LLMs for English within the same timeframe. While primarily trained for English, these models’ vocabulary encompasses tokens for other languages as well. Given their substantial pre-training data, which may include samples from other languages, Llama models can also exhibit promising performance in multilingual scenarios. We engage the 7B parameter model for both Llama 2 and Llama 3 in our evaluation process.
@@
  • MEDITRON [6], released in November 2023 is an open-source monolingual biomedical LLM leveraging extra 45B English tokens for further pre-training the general LLM Llama 2. It has two scaling versions, i.e., 7B and 70B and for fair comparison with others, we mainly adopt the 7B version.
  • BioMistral [23], released in February 2024, is an open-source multilingual biomedical LLM based on Mistral. It is concurrent with ours and is also targeted at the multilingual biomedical domain. We compare with it as a strong baseline.
  • Gemma [41], released in March 2024, is an open-source monolingual LLM developed by Google DeepMind, targeting in English. It demonstrates strong performance across academic benchmarks for language understanding, reasoning, and safety. It has two versions, i.e., 2B and 7B scales. Similarly, for a fair comparison, we adopt the 7B one herein.
We have concluded more detailed information for each model in Supplementary Material B.
Metrics and Human Rating. In this part we will introduce the evaluation metrics and human rating criteria used in our work. To evaluate the performance of LLMs, we employ two metrics: Accuracy and Rationale Similarity. Measuring Accuracy is straightforward, as the LLM can generate outputs following a specific template. However, assessing Rationale Similarity presents a more complex challenge, typical within the NLP domain. Initially, we applied three classical text similarity methods, BLEU [36] and ROUGE [26], and Bert-score [52].
@@
BERT-score utilizes the contextual embedding from a pre-trained BERT to capture high-level semantic features, calculating the similarity between reference and candidate texts through cosine similarity. The Recall is calculated as follows:
@@
Relative Rating scores aim to rank the output based on relative comparison. In detail, we select 6 representative models and sample 50 cases for each language. In human rating, for each case, question, options, right answer and rationale generated by each model are present to annotators, along with the reference rationale. The annotators are asked to rank the model-generated rationales based on the following three evaluation criteria:
@@
4.4 English Benchmark Evaluation
Here, we describe how we compare the performance of our model on English with other existing models.
In assessing the capabilities of large language models in medical field, we utilize 4 widely recognized multiple-choice question-answering benchmarks, as follows :
  • MedQA [17] is the same as we introduced in MMedBench. It is a widely used and highly credible benchmark for assessing the medical ability of models. Thus we re-use it in English-wise evaluation.
  • PubMedQA [17] is an English question-answering medical dataset based on PubMed abstracts. The task of PubMedQA is to answer research questions with yes/no/maybe, which can also be viewed as a close-domain multiple-choice question-answering problem. Officially, it is split into three subsets: 1K manually labeled pairs (PQA-L), 61.2K unlabeled pairs (PQA-U), and 211.3K artificially generated pairs (PQA-A). Following former existing works [9], we also adopt PQA-L as the test set so that our results can be compared with others directly.
  • MedMCQA [35] is a large-scale English multiple-choice question-answering samples. MedMCQA has more than 194k high-quality AIIMS & NEET PG entrance exam multiple-choice questions covering 2.4k healthcare topics and 21 medical subjects are collected with an average token length of 12.77 and high topical diversity. The official train split contains 182,822 questions, and the test split contains 4183 questions. Each question has 4 choices. We adopt the official test split to evaluate our model.
  • MMLU-Medicine [12] is an English comprehensive large-scale exam questions spanning 57 subjects, aiming to assess the ability of language models across different domains. Following MedPALM 2 [39], we adopt the 6 subjects related to medicine, i.e., anatomy (An), clinical knowledge (CK), college biology (CB), college medicine (CM), professional medicine (PM) and medical genetics (MG), featuring 1,089 questions. We adopt the official split of MMLU for testing.
In English, LLMs may use the mixture of supervised data to further align the model with human semantic instructions after pre-training, commonly referred to as instruction tuning [7, 27, 47]. This setup is similar to fine-tuning, but the difference is that instruction tuning often involves designing semantic instructions to aggregate a large number of tasks, rather than just considering a few downstream tasks to be tested. Different LLMs may use different dataset collections for instruction tuning. Thus in English benchmarks it is hard to control the data, like that we perform in the fine-tuning setting on MMedBench, instead, models are compared directly on the unexposed testing set regardless of what tuning data they used. In our case, to enable a fair comparison with existing models, we incorporate an off-the-shelf English instruction fine-tuning dataset (from PMC-LLaMA [48]) into MMed-Llama 3 finetuning.
@@ 4.5 Implementation Details
Auto-regressive Training. During further auto-regressive training on MMedC, our optimization objective aligns with that of the auto-regressive generation task. For data processing, we segment the text into chunks, each comprising 2048 tokens, with an overlapping margin of 512 tokens. Throughout the training, we maintain a maximum context length of 2048 tokens. Owing to the model’s extensive parameter count, which precludes fitting on a single GPU, we employ the Fully Sharded Data Parallel (FSDP) strategy to distribute the model across multiple GPUs. Additionally, we utilize the BF16 data type and gradient checkpointing techniques to optimize memory usage. For InternLM, we establish a global batch size of 512 and a learning rate of 2e-5. In the case of BLOOM, we set the global batch size to 512 and a learning rate of 8e-6. We pre-train both models on eight A100 GPUs, adapting gradient accumulation steps to sustain such a large global batch size. We train the 7B model for 20k iterations, which takes about 20 Days.
@@ 5 Conclusion
In this paper, we describe an automatic pipeline for building up a new multilingual medical corpus together with a new benchmark. Specifically, we developed MMedC, a large-scale medical corpus with 25.5B tokens covering six main languages, and MMedBench, a comprehensive benchmark encompassing 6 primary languages. To monitor the development progress of medical multilingual LLMs, we have evaluated eleven existing LLMs on multi-choice question-answering and rationale generation abilities under various settings. Experimentally, we demonstrate the effectiveness of further training on MMedC, significantly filling the gap in adapting advanced general multilingual LLMs into complex and professional medical domain. As a result, we open-source the MMed-Llama 3, which to our knowledge, denotes the first, and strongest multilingual language model for medicine, also demonstrating impressive performance in various English benchmarks.
6 Data Availability
Source data are provided in this paper. MMedC Dataset is released in https://huggingface.co/datasets/Henrychur/MMedC with CC BY-NC-SA license. MMedBench Benchmark Data is released in https://huggingface.co/datasets/Henrychur/MMedBench with CC BY-NC-SA license. MMedBench Leaderboard can be found in https://henrychur.github.io/MultilingualMedQA/.
7 Code Availability
source code of this paper is released in https://github.com/MAGIC-AI4Med/MMedLM with CC BY-SA license. MMedLM model weights can be found in https://huggingface.co/Henrychur/MMedLM and MMedLM 2 model weights is in https://huggingface.co/Henrychur/MMedLM2 all with Apache-2.0 license [^1] https://www.apache.org/licenses/LICENSE-2.0, the same as the base InternLM-series. MMed-Llama 3 can be found in https://huggingface.co/Henrychur/MMed-Llama-3-8B and the English-wise instruction-tuned model can be found in https://huggingface.co/Henrychur/MMed-Llama-3-8B-EnIns with the Llama 3 community license [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B/blob/main/LICENSE
@@ A.0.1 Zero-shot prompt
For models such as GPT-3.5 and GPT-4, we employ specific instructions to steer the model’s response to questions. To prompt the model for a straightforward correct answer, we use the following directive: {mdframed}[backgroundcolor=gray!20] You’re a {language} doctor, make a choice based on the question and options. You need to answer the letter of the option instead of answering the entire option or anything else. Options may not be unique. When we aim for the model to provide not just the correct answer but also its corresponding rationale, we utilize a different instruction: {mdframed}[backgroundcolor=gray!20] You’re a {language} doctor, make a choice based on the question and options in {language}. You should solve this step-by-step. You must first give the reason in {language} for your choice ends with ’[End]’. Then you must give the answer’s letter directly again. The template is like ’Reason:… [End] Answer: A, B’